College Physics Problem. Please show all work

An object with a mass of 4.55 kg starts from rest at point o and moves in the xy plane under the influence of the force F = (7xǐ + 2 N/m, where x and y are in meters y (m) (m Determine the work done on the object by the force as it moves along the paths listed below. (Include the sign of the values in your answers.) (a) path oC (b) path OBC (c) round trip by path OBCO (d) From the above, we may conclude that the given force is which of the following? O is not a conservative force O not enough information to tell O is a conservative force
Answer #1

mass of object {mp=4.55kg. Force (F) = 72h + 2y } Xao = 56J _123 @) along hath 06 [60-) —- (4,0)]. now w=f7-05 = J 1739 +ay 7